### v1

# HPSS: Heuristic Prompting Strategy Search for LLM Evaluators

Abstract
Since the adoption of large language models (LLMs) for text evaluation has become increasingly prevalent in the field of natural language processing (NLP), a series of existing works attempt to optimize the prompts for LLM evaluators to improve their alignment with human judgment. However, their efforts are limited to optimizing individual factors of evaluation prompts, such as evaluation criteria or output formats, neglecting the combinatorial impact of multiple factors, which leads to insufficient optimization of the evaluation pipeline. Nevertheless, identifying well-behaved prompting strategies for adjusting multiple factors requires extensive enumeration. To this end, we comprehensively integrate 8 key factors for evaluation prompts and propose a novel automatic prompting strategy optimization method called Heuristic Prompting Strategy Search (HPSS). Inspired by the genetic algorithm, HPSS conducts an iterative search to find well-behaved prompting strategies for LLM evaluators. A heuristic function is employed to guide the search process, enhancing the performance of our algorithm. Extensive experiments across four evaluation tasks demonstrate the effectiveness of HPSS, consistently outperforming both human-designed evaluation prompts and existing automatic prompt optimization methods.
HPSS: Heuristic Prompting Strategy Search for LLM Evaluators
Bosi Wen1,†  Pei Ke2 Yufei Sun4 Cunxiang Wang3,4 Xiaotao Gu3 Jinfeng Zhou1 Jie Tang4 Hongning Wang1  Minlie Huang1,‡ 1The Conversational Artificial Intelligence (CoAI) Group, Tsinghua University 2University of Electronic Science and Technology of China  3Zhipu AI 4The Knowledge Engineering Group (KEG), Tsinghua University wbs23@mails.tsinghua.edu.cn, aihuang@tsinghua.edu.cn
Work done when this author interned at Zhipu AI.Corresponding author
1 Introduction
Evaluation is a long-standing critical and challenging task in NLP Celikyilmaz et al. (2020); Chang et al. (2024). Recently, with the advent of advanced large language models (LLMs) such as GPT-4, LLM-based evaluators have been widely used for various natural language generation tasks Wang et al. (2023a); Liu et al. (2023); Zheng et al. (2023); Hu et al. (2024). Leveraging the LLMs’ strong language understanding and instruction-following capabilities, the evaluation task can be addressed via an evaluation prompt that includes a task description, evaluation rules, and the text to be assessed Chen et al. (2023); Ke et al. (2023). Various studies have reported that LLMs can provide accurate and interpretable evaluation results of text quality on the specified aspect, offering stronger generalization ability compared with previous evaluation methods Papineni et al. (2002); Lin (2004); Zhang et al. (2020); Zhong et al. (2022).
Figure 1: An example of prompting strategy and its corresponding evaluation prompt for assessing the coherence of text summarization, where some factors are highlighted and others are not shown.
Despite the advantages mentioned above, LLM-based evaluation methods are still imperfect in aligning with human judgments Wang et al. (2023a); Liu et al. (2024b), which cannot fully serve as an alternative to human evaluation as the gold standard. An emerging line of research attempts to optimize evaluation prompts for LLM evaluators to improve their alignment with human judgments from multiple aspects, including optimizing evaluation criteria Liu et al. (2024c, d), modifying output formats Chiang and Lee (2023); Chu et al. (2024), and providing in-context examples Kim et al. (2023); Jain et al. (2023); Huang et al. (2024).
However, we argue that these methods are only limited to optimizing a single factor of evaluation prompts (e.g., evaluation criteria or output formats) but neglect the prompting strategy for adjusting multiple factors (Kim et al., 2023). Given that the evaluation prompts comprise multiple components Gao et al. (2024), each encompassing various factors that simultaneously affect the performance of LLM evaluators, it is necessary to collectively adjust these factors to fully harness its evaluation capability. Nevertheless, due to the vast state space involved in making these adjustments, extensive enumeration is required to identify well-behaved prompting strategies, highlighting the need for effective automatic optimization methods.
To this end, we comprehensively integrate a series of prompting strategies for LLM evaluators from previous works, select 8 key factors, and propose the Heuristic Prompting Strategy Search (HPSS) algorithm to automatically optimize prompting strategies for LLM evaluators tailored to specific evaluation aspects. Inspired by the genetic algorithm Mitchell (1980), HPSS gradually mutates the previous prompting strategies and selects candidate strategies based on their performance on a validation dataset. Leveraging the limited number of values within each factor, HPSS introduced a heuristic function to guide the search process, assigning higher exploration probability to more promising values, which enhances the effectiveness of mutation and leads to better performance compared to random search. We optimize the prompting strategy for two different LLM evaluators (i.e., GPT-4o-mini OpenAI (2023) and Qwen2.5-14B-Instruct Yang et al. (2024)) across four evaluation tasks for natural language generation (NLG). Experiment results show that HPSS achieves substantial improvement over the baseline of human-designed evaluation prompts and existing automatic prompt optimization methods. Compared to the commonly-used evaluation prompt from MT-Bench Zheng et al. (2023), HPSS can achieve an average 29.4% relative performance improvement with the same generation times. Furthermore, compared to other manually designed prompting strategies such as G-Eval Liu et al. (2023) and CloserLook Chiang and Lee (2023), HPSS can still achieve significantly better performance with only 5% of the generation times. Our contributions can be summarized as follows:
  • To the best of our knowledge, we present the first discussion of automatic prompting strategy optimization for LLM evaluators, demonstrating that appropriate prompting strategies can significantly enhance the performance of LLM evaluators.
  • We propose HPSS, a novel prompting strategy optimization algorithm via iterative search. Furthermore, a heuristic function is introduced to estimate the prospect of each value and enhance the effectiveness of mutation.
  • We validate the effectiveness of HPSS across four evaluation tasks, yielding consistently better performance compared to both human-designed evaluation prompts and existing automatic prompt optimization methods.
2 Related Work
2.1 Prompt Design for LLM Evaluators
With the emergence of LLMs, utilizing LLMs as evaluators to assess the quality of given texts has gradually become prevalent Wang et al. (2023a); Chen et al. (2023); Ke et al. (2023) due to its advantages in flexibility, interpretability, and generalization. To enhance their performance, recent researches attempt to optimize the evaluation prompts for LLM evaluators from multiple aspects Stureborg et al. (2024); Kim et al. (2023); Jain et al. (2023); Murugadoss et al. (2024); He et al. (2024); Pereira and Lotufo (2024). Specifically, G-Eval Liu et al. (2023) requires LLMs to generate evaluation steps first and refer to these steps to finish the evaluation task. LLMBar Zeng et al. (2024) explores the effects of self-generated reference answers and sample-specific metrics on text evaluation. HD-Eval Liu et al. (2024d) attempts to align LLM evaluators with humans through hierarchical evaluation criteria decomposition. However, these works are constrained to optimize an individual factor for evaluation prompts, leading to insufficient optimization. In contrast, our work focuses on adjusting multiple factors and searching for effective prompting strategies to fully stimulate the potential of LLM evaluators.
2.2 Prompt Optimization
Factor Definition Common Usage Selection Range
Scoring
Scale
The scoring range for text evaluation. Various scoring scales are used in previous work, such as 1-3 Gopalakrishnan et al. (2019); Lin and Chen (2023), 1-5 Fabbri et al. (2021); Chhun et al. (2022), 1-10 Liu et al. (2024a) and 1-100 Stureborg et al. (2024). 1-3 1-5 1-10 1-50 1-100
In-Context
Example
Human evaluation examples Previous works try to construct in-context examples through random or stratified sampling based on human ratings to enhance the performance of LLM evaluators Kim et al. (2023); Huang et al. (2024); Jain et al. (2023). The number of examples usually ranges from 1 to 4. 0 examples 3 examples 5 examples 10 examples
Evaluation
Criteria
The general definition of the evaluation aspect and the scoring standards of the quality of a text. The majority of previous works use human-written criteria. Some recent works also prompt LLM evaluators to generate criteria by themselves Kotonya et al. (2023) or even use no criteria Murugadoss et al. (2024). No Criteria Human-Written Criteria Self-Generated Criteria
Reference The reference answer for the task in the sample to evaluate. In addition to using no reference Liu et al. (2023), previous work also utilizes the following types of references: (1) Self-Generated Reference: prompting the evaluator to generate a reference independently and then feed it into $\mathcal{T}$ Zeng et al. (2024), (2) Dialectic: prompting the evaluator to generate a reference along with the final rating He et al. (2024) No Reference Self-Generated Reference Dialectic
Chain-of-Thought
Request for the LLM evaluator to generate evaluation explanations Previous work mainly utilizes following CoT formats: (1) No CoT: generate the rating without an explanation Liu et al. (2023), (2) Prefix CoT: provide an explanation first and then give the rating Zheng et al. (2023); Chiang and Lee (2023), and (3) Suffix CoT: provide the rating first, followed by an explanation Chiang and Lee (2023) No CoT Prefix CoT Suffix CoT
AutoCoT Self-generated evaluation steps for specific evaluation aspect G-Eval Liu et al. (2023), OP-I-Prompt Siledar et al. (2024) No AutoCoT AutoCoT
Metrics Self-generated sample-specific metrics for good answer LLMBar Zeng et al. (2024), Check-Eval Pereira and Lotufo (2024) No Metrics Metrics
Order The placement order of each component in $\mathcal{T}$ For three main components of $\mathcal{T}$, i.e., Task Description (TD), Evaluation Rule (ER), and Input Content (IC), there are two commonly-used placement orders: TD $\rightarrow$ ER $\rightarrow$ IC Zheng et al. (2023); Liu et al. (2023) and TD $\rightarrow$ IC $\rightarrow$ ER Liu et al. (2024c). TD $\rightarrow$ ER $\rightarrow$ IC TD $\rightarrow$ IC $\rightarrow$ ER ER $\rightarrow$ TD $\rightarrow$ IC ER $\rightarrow$ IC $\rightarrow$ TD IC $\rightarrow$ TD $\rightarrow$ ER IC $\rightarrow$ ER $\rightarrow$ TD
Table 1: Factors of evaluation prompts, including their definition, common usage in previous works, and selection range considered in our work. The definitions of each component in the last row are described in Section 3.2.
Although LLMs have demonstrated impressive performance across various NLP tasks Zhao et al. (2023), researchers point out that their abilities are highly dependent on sophisticated prompt design Pryzant et al. (2023); Leidinger et al. (2023); Raina et al. (2024). Considering that manual prompt engineering is time-consuming, many works attempt to optimize prompt automatically. GPS Xu et al. (2022) and GRIPS Prasad et al. (2023) employ a genetic algorithm, using rule or generative language models to mutate the initial prompt. PromptBreeder Fernando et al. (2023) and EvoPrompt Guo et al. (2023) utilize LLMs for mutation and crossover operations in evolutionary searches. OPRO Yang et al. (2023) demonstrates the potential of LLMs to perform iteratively prompt optimization based on search history. Furthermore, Hsieh et al. (2024) extends prompt optimization to long prompts containing multiple sentences. They utilized Lin-UCB to guide the selection of sentences and employed in-context learning for LLM-based mutation. For comparison, our work shifts the focus from word or sentence-level prompt optimization to a higher-level strategy that adjusts multiple factors of evaluation prompts. Additionally, we introduce a heuristic function to enhance the effectiveness of mutation based on the characteristics of evaluation prompting strategies.
3 Methodology
Figure 2: Overview of our HPSS algorithm. By perturbing the value of each factor in the prompting strategy from MT-Bench, HPSS gets the initial strategy population and the advantage estimations of values. Subsequently, new strategies are iteratively searched based on the guidance of the value advantage. The performance of new strategies is used to update the top-$k$ strategy population and the advantage estimation.
3.1 Problem Formulation
NLG evaluation tasks typically require LLM evaluators to generate the corresponding judgment given the input sample $d$ and evaluation aspect $a$. The input sample may include the query and model-generated text. Meanwhile, the judgment takes various forms, including the rating score of the generated text in pointwise grading and the preference label between two generated texts in pairwise comparison. To guide LLM evaluators to assess the quality of the input sample $d$ within the aspect $a$, a prompt template $\mathcal{T}$ is needed to provide sufficient clarifications of the task. In practice, $d$ and $a$ will be fed into $\mathcal{T}$ to constitute an evaluation instruction, and the LLM evaluator will follow this instruction to generate the final judgment.
Since $\mathcal{T}$ contains various factors, there are multiple prompting strategies for adjusting these factors. An example of prompting strategy is illustrated in Figure 1. Formally, assume that $\mathcal{T}$ has $n$ factors $(F_{1},F_{2},\ldots,F_{n})$, each factor $F_{i}$ has $m_{i}$ possible values $(f_{i1},f_{i2},\ldots,f_{im_{i}})$. Thus, $\mathcal{T}$ can be determined by $\boldsymbol{F}=(F_{1},F_{2},\ldots,F_{n})$. For a dataset $D=\{d_{i}\}_{i=1}^{|D|}$ and an evaluation aspect $a$, our objective is to search for the prompting strategy $\mathcal{T}^{o}$ to maximize the performance of LLM evaluators:
$\mathcal{T}^{o}=\arg\max_{\boldsymbol{F}}r_{D,a}(\mathcal{T}_{\boldsymbol{F}})$ (1)
where $r_{D,a}(\mathcal{T})$ represents the performance of $\mathcal{T}$ for the aspect $a$ on the dataset $D$, which is typically measured by the correlation between the evaluation results of LLM evaluators using $\mathcal{T}$ and human judgments. To aid the search for the optimal prompting strategy, we will use a validation dataset with human judgments to evaluate the performance of $\mathcal{T}$. Without loss of generality, we simplify $r_{D,a}(\mathcal{T})$ as $r(\mathcal{T})$ to denote the performance of $\mathcal{T}$ for a specific aspect on the validation dataset.
3.2 Factors of Evaluation Prompt
Considering a series of prompting strategies from previous works, we select 8 key factors for the prompting strategy search, as described in Table 1. Detailed prompts are provided in Appendix A.
The first 7 factors belong to 3 main components of the evaluation prompt template $\mathcal{T}$: (1) Task Description (TD), which encompasses the definition of the evaluation task and output format (including Scoring Scale and Chain-of-Thought), (2) Evaluation Rules (ER), which encompass the standard and steps for evaluation (including Evaluation Criteria and AutoCoT), and (3) Input Content (IC), which encompasses the input sample and sample-specific auxiliary information (including In-Context Example, Reference, and Metrics). Previous works adopt different orders for these 3 components. For instance, MT-Bench Zheng et al. (2023) uses the order TD $\rightarrow$ ER $\rightarrow$ IC, while AutoCalibrate Liu et al. (2024c) uses TD $\rightarrow$ IC $\rightarrow$ ER. We consider the order of these 3 components as the last factor and all 6 possible orders as the selection range. The overall size of the search space is 12,960. We also conduct a preliminary experiment to explore the effect of these factors on Appendix B. The results show that these factors significantly influence the performance of LLM evaluators, yet some findings diverge from previous works.
3.3 Heuristic Prompting Strategy Search
Inspired by the Genetic Algorithms (GA) Mitchell (1980), we propose Heuristic Prompting Strategy Search (HPSS) to effectively search prompting strategies to help LLM evaluators achieve better alignment with human judgment. The framework is shown in Figure 2. HPSS maintains a population of $k$ top-performing prompting strategies and mutates each strategy in this population to create new strategies. Their performance is measured by the human correlation between LLM evaluators and human judgment on the validation dataset. Different from traditional GA that randomly selects values for mutation with equal probability, HPSS leverages the limited number of values within each factor and calculates the expected correlation metric advantage of each value over the random selection of the corresponding factor. The advantage of value $f_{ij}$ for factor $F_{i}$ denotes as $A_{ij}$. Then, this advantage score serves as a heuristic function to guide mutation. This mechanism assigns higher selection probabilities to more promising values, reducing the cost caused by the blindly random search in GA. HPSS mainly contains two steps:
Model Method Summeval Topical-Chat
Coherence Consistency Fluency Relevance Avg. Coherence Engagingness Groundedness Naturalness Avg.
- BLEU-4 0.021 0.043 0.066 0.154 0.071 0.403 0.447 0.325 0.352 0.382
- BertScore 0.228 0.194 0.225 0.370 0.254 0.390 0.438 0.351 0.335 0.379
- UniEval 0.623 0.469 0.455 0.426 0.493 0.589 0.597 0.553 0.487 0.557
GPT-4o-mini MT-Bench 0.481 0.482 0.319 0.445 0.432 0.546 0.635 0.530 0.536 0.562
G-Eval 0.517 0.532 0.315 0.517 0.470 0.576 0.669 0.634 0.579 0.615
CloserLook 0.581 0.506 0.431 0.532 0.513 0.551 0.705 0.630 0.655 0.635
CloserLook + ICL 0.540 0.538 0.469 0.522 0.517 0.540 0.676 0.619 0.564 0.600
CloserLook + ICL 0.554 0.514 0.522 0.539 0.532 0.592 0.740 0.652 0.655 0.660
APE 0.560 0.516 0.401 0.514 0.498 0.542 0.668 0.705 0.562 0.619
OPRO 0.529 0.495 0.493 0.532 0.512 0.641 0.727 0.792 0.618 0.695
Greedy 0.577 0.516 0.477 0.554 0.531 0.614 0.715 0.668 0.675 0.668
Stepwise-Greedy 0.545 0.466 0.494 0.552 0.514 0.619 0.713 0.677 0.630 0.662
HPSS (Ours) 0.601 0.567 0.525 0.561 0.564 0.661 0.743 0.866 0.690 0.740
Qwen2.5-14B- Instruct MT-Bench 0.449 0.460 0.455 0.451 0.454 0.394 0.635 0.517 0.497 0.511
G-Eval 0.569 0.544 0.467 0.515 0.524 0.490 0.673 0.505 0.535 0.550
CloserLook 0.602 0.502 0.476 0.530 0.528 0.510 0.706 0.646 0.555 0.604
CloserLook + ICL 0.478 0.587 0.355 0.480 0.475 0.436 0.675 0.642 0.580 0.583
CloserLook + ICL 0.554 0.504 0.476 0.516 0.512 0.520 0.712 0.715 0.595 0.636
APE 0.489(0.011) 0.494(0.019) 0.451(0.021) 0.487(0.025) 0.480 0.466(0.025) 0.670(0.034) 0.567(0.039) 0.591(0.037) 0.574
OPRO 0.479(0.041) 0.564(0.003) 0.481(0.003) 0.498(0.010) 0.506 0.502(0.071) 0.737(0.000) 0.692(0.000) 0.583(0.000) 0.629
Greedy 0.528(0.023) 0.530(0.028) 0.489(0.007) 0.527(0.021) 0.519 0.511(0.002) 0.728(0.030) 0.776(0.000) 0.543(0.034) 0.640
Stepwise-Greedy 0.553(0.000) 0.516(0.000) 0.491(0.000) 0.528(0.000) 0.522 0.586(0.000) 0.721(0.000) 0.635(0.000) 0.601(0.000) 0.636
HPSS (Ours) 0.559(0.000) 0.625(0.007) 0.515(0.008) 0.542(0.006) 0.560 0.598(0.030) 0.760(0.014) 0.795(0.016) 0.635(0.017) 0.697
- w/o exploitation 0.543(0.012) 0.626(0.008) 0.511(0.009) 0.538(0.003) 0.555 0.537(0.038) 0.741(0.015) 0.782(0.009) 0.596(0.021) 0.664
- w/o exploration term 0.550(0.012) 0.623(0.012) 0.508(0.005) 0.527(0.008) 0.552 0.555(0.026) 0.760(0.013) 0.789(0.018) 0.611(0.000) 0.679
- w/o all 0.525(0.010) 0.617(0.004) 0.499(0.005) 0.536(0.002) 0.544 0.516(0.026) 0.745(0.021) 0.769(0.005) 0.568(0.002) 0.650
Table 2: Summary-level Spearman correlations of different aspects on Summeval and dataset-level Spearman correlations on Topical-Chat. †  indicates that the corresponding method employs 20 generations with self-consistency.
Initialization.
This step aims to get the initial strategy population and advantage estimation of values. Starting with a commonly-used reference-free pointwise grading prompting strategy from MT-Bench, we modify the selection value for each factor separately to construct multiple strategies. After that, we calculate the initial advantage estimation of each value based on their performance and select $k$ top-performing strategies for the subsequent iterative search step.
Iterative Search.
In each iteration, HPSS mutates each prompting strategy $\mathcal{T}^{c}$ within the population for $g$ times to generate new strategies, and then update the top-$k$ strategy population based on the performance of these new strategies. To enhance the effectiveness of mutation, HPSS takes overall advantage of all factor values as the performance estimation of one strategy and allocates more search resources to the strategies with higher overall advantages. Specifically, we utilize two types of mutations: exploration and exploitation, which have the probability $\rho$ and $1-\rho$ to be chosen for each mutation, respectively.
(1) Exploration: Exploration selects the strategies adjacent to $\mathcal{T}^{c}$ (i.e., the strategies set $T_{adj}$ which can be obtained by modifying a single factor of $\mathcal{T}^{c}$) and calculates the disparity between the overall advantages of these strategies and $\mathcal{T}^{c}$, employing a temperature-controlled softmax function to determine the exploration probability. Furthermore, in the preliminary experiment, we notice that inaccurate advantage estimation during Initialization stage may lead to insufficient exploration of well-performing values. To address this, motivated by UCB Sampling Lai and Robbins (1985), we include an additional exploration term in softmax sampling to incentivize a more comprehensive exploration of all the values. Formally, assume the value for $F_{i}$ within $\mathcal{T}^{c}$ is $f_{ic_{i}}$, and the strategy $\mathcal{T}$ in $T_{adj}$ modifies the value for $F_{i}$ from $f_{ic_{i}}$ to $f_{ij}$, then the exploration probability is computed as follows:
$B_{\mathcal{T}}=A_{ij}-A_{ic_{i}}+\lambda\sqrt{\ln(t)/{M_{ij}}}$ (2)
$P(\mathcal{T})=\frac{\exp(B_{\mathcal{T}}/\tau)}{\sum_{\mathcal{T}^{\prime}\in T_{adj}}\exp(B_{\mathcal{T}^{\prime}}/\tau)}$ (3)
where $t$ is the search step, and $M_{ij}$ is the appearance count of value $f_{ij}$ during search process.
After exploring a new prompting strategy $\mathcal{T}^{new}$, we evaluate its performance on the validation dataset and calculate the performance gain brought by the modified value. After that, a moving average with previous results is applied to update the value’s advantage estimation. The advantages of the values that belong to the same factor as the modified value will then be normalized to have a zero mean, so as to satisfy the property that the expectation sums to zero. Formally, the advantage $A$ will be updated as follows, where $N_{ij}$ is the exploration count of the value $f_{ij}$:
$A_{ij}\leftarrow\frac{A_{ij}\cdot N_{ij}+[r(\mathcal{T}^{new})-(r(\mathcal{T}^{c})-A_{ic_{i}})]}{N_{ij}+1}$ (4)
$A_{ik}\leftarrow A_{ik}-avg(A_{i1},A_{i2},\cdots,A_{i{m_{i}}}),$ (5)
$\hskip 50.00008ptk=1,2,\cdots,m_{i}$
(2) Exploitation: Since the search scope of exploration is constrained to the vicinity of explored prompting strategies, HPSS also employs the mutation type exploitation to expand the search scope and avoid getting stuck in a local optimum. This method selects the unexplored strategy $\mathcal{T}^{max}$ with the highest overall advantage:
$\mathcal{T}^{max}=\mathop{\arg\max}\limits_{\boldsymbol{F}}\sum_{i=1}^{n}A_{F_{i}}$ (6)
After reaching the maximum search step, HPSS returns the best-performing strategy on the validation dataset as the final result. Detailed implementation of these two steps is presented in Appendix H.
4 Experiments
4.1 Experimental Setup
Tasks and Datasets
We evaluate HPSS on four pointwise grading evaluation tasks: Summeval Fabbri et al. (2021) for text summarization, Topical-Chat Gopalakrishnan et al. (2019) for dialogue generation, SFHOT / SFRES Wen et al. (2015) for data-to-text generation, and HANNA Chhun et al. (2022) for story generation. The Spearman ($\rho$) correlation coefficient between human judgments and LLM evaluations is adopted as the performance metric. We also validate HPSS on pairwise comparison benchmarks MT-Bench Zheng et al. (2023), AUTO-J Li et al. (2024), and LLMBar Zeng et al. (2024). The results are provided in Appendix I. More details about benchmarks and metrics can be found in Appendix D and E. Following HD-Eval Liu et al. (2024d), for each dataset, a 50% proportion is held out for testing, while the rest is applied for validation.
Baselines
We compare HPSS with three types of baselines: (1) Non-LLM Evaluators: This category includes BLEU-4 Papineni et al. (2002), BertScore Zhang et al. (2020), and UniEval Zhong et al. (2022). (2) Human-Designed LLM Evaluators: The prompting strategy from MT-Bench Zheng et al. (2023) stands as the starting point of searching. G-Eval Liu et al. (2023) integrates AutoCoT to enhance the performance of LLM evaluators. Chiang and Lee (2023) explores various evaluation schemes. We use its best setting analyze-rate, denoted as CloserLook, and attempt to further improve its performance by providing human evaluation examples (CloserLook + ICL). (3) Automatic Prompt Optimization for LLM Evaluators: We implement 4 strong prompt optimization baselines, including APE Zhou et al. (2023b), OPRO Yang et al. (2023), Greedy Prasad et al. (2023); Zhou et al. (2023a) and Stepwise-Greedy. Greedy is a widely used iterative search algorithm, where random perturbations are applied to the current strategy to generate multiple candidates at each iteration, and the best-performing one is retained for the next iteration. Stepwise-Greedy sequentially optimizes each factor, choosing the best-performing value in each step. More details can be found in Appendix F.
Model Method SFHOT SFRES HANNA
Informativeness Naturalness Avg. Informativeness Naturalness Avg. Relevance Coherence Empathy Surprise Engagement Complexity Avg.
- BLEU-4 0.046 0.029 0.038 0.227 0.142 0.185 0.308 0.313 0.307 0.244 0.329 0.374 0.313
- BertScore 0.103 0.076 0.090 0.213 0.150 0.182 0.298 0.383 0.403 0.331 0.387 0.436 0.373
- UniEval 0.237 0.312 0.275 0.211 0.340 0.276 - - - - - - -
GPT-4o-mini MT-Bench 0.307 0.385 0.346 0.253 0.417 0.335 0.499 0.457 0.475 0.365 0.490 0.500 0.464
G-Eval 0.288 0.421 0.354 0.313 0.420 0.367 0.507 0.564 0.446 0.344 0.500 0.540 0.484
CloserLook 0.325 0.439 0.382 0.312 0.395 0.353 0.586 0.570 0.536 0.443 0.559 0.615 0.552
CloserLook + ICL 0.327 0.445 0.386 0.295 0.379 0.337 0.528 0.518 0.493 0.384 0.552 0.575 0.508
CloserLook + ICL 0.358 0.447 0.402 0.292 0.402 0.347 0.566 0.590 0.539 0.425 0.585 0.621 0.555
APE 0.289 0.427 0.358 0.272 0.373 0.323 0.524 0.494 0.507 0.321 0.427 0.499 0.462
OPRO 0.344 0.403 0.374 0.328 0.378 0.353 0.508 0.550 0.469 0.388 0.536 0.583 0.506
Greedy 0.371 0.435 0.403 0.284 0.427 0.356 0.548 0.584 0.438 0.433 0.537 0.566 0.518
Stepwise-Greedy 0.368 0.452 0.410 0.292 0.382 0.337 0.566 0.517 0.522 0.424 0.537 0.521 0.515
HPSS (Ours) 0.395 0.466 0.431 0.370 0.439 0.405 0.548 0.594 0.535 0.459 0.542 0.591 0.545
Qwen2.5-14B- Instruct MT-Bench 0.281 0.361 0.321 0.181 0.311 0.241 0.447 0.459 0.387 0.286 0.430 0.427 0.406
G-Eval 0.262 0.418 0.340 0.308 0.455 0.382 0.514 0.467 0.444 0.326 0.391 0.504 0.441
CloserLook 0.266 0.419 0.342 0.247 0.445 0.346 0.537 0.469 0.493 0.369 0.436 0.539 0.474
CloserLook + ICL 0.245 0.411 0.328 0.281 0.396 0.339 0.435 0.440 0.437 0.325 0.476 0.461 0.429
CloserLook + ICL 0.241 0.419 0.330 0.336 0.406 0.371 0.560 0.532 0.501 0.373 0.506 0.538 0.502
APE 0.310(0.016) 0.377(0.025) 0.344 0.293(0.033) 0.426(0.009) 0.360 0.509(0.026) 0.476(0.004) 0.456(0.019) 0.321(0.016) 0.417(0.018) 0.450(0.021) 0.452
OPRO 0.301(0.000) 0.358(0.000) 0.330 0.326(0.025) 0.388(0.000) 0.357 0.580(0.004) 0.517(0.015) 0.436(0.000) 0.318(0.030) 0.517(0.009) 0.550(0.029) 0.486
Greedy 0.296(0.017) 0.381(0.016) 0.339 0.294(0.001) 0.425(0.023) 0.360 0.552(0.011) 0.493(0.010) 0.468(0.001) 0.325(0.034) 0.495(0.011) 0.541(0.001) 0.479
Stepwise-Greedy 0.291(0.000) 0.345(0.000) 0.318 0.333(0.000) 0.359(0.000) 0.346 0.506(0.000) 0.462(0.000) 0.459(0.000) 0.383(0.000) 0.524(0.000) 0.590(0.000) 0.487
HPSS (Ours) 0.321(0.000) 0.411(0.000) 0.366 0.351(0.005) 0.444(0.002) 0.398 0.586(0.004) 0.538(0.018) 0.463(0.005) 0.386(0.020) 0.545(0.001) 0.594(0.007) 0.519
- w/o exploitation 0.295(0.008) 0.398(0.010) 0.347 0.346(0.006) 0.425(0.013) 0.386 0.544(0.010) 0.517(0.003) 0.481(0.017) 0.370(0.007) 0.514(0.010) 0.539(0.004) 0.494
- w/o exploration term 0.296(0.006) 0.403(0.012) 0.350 0.347(0.008) 0.433(0.011) 0.390 0.561(0.027) 0.516(0.003) 0.460(0.000) 0.378(0.012) 0.521(0.005) 0.571(0.011) 0.501
- w/o all 0.284(0.022) 0.387(0.014) 0.336 0.348(0.009) 0.413(0.002) 0.381 0.514(0.012) 0.524(0.007) 0.467(0.006) 0.359(0.002) 0.483(0.008) 0.521(0.009) 0.478
Table 3: Dataset-level Spearman correlations of different aspects on SFHOT, SFRES, and HANNA. †  indicates that the corresponding method employs 20 generations with self-consistency.
Models and Configurations
We choose the representative closed-source model GPT-4o-mini OpenAI (2023) and open-source model Qwen2.5-14B-Instruct Yang et al. (2024) as evaluation models. For G-Eval and CloserLook, we use their default settings (with 20 generation times and the decoding temperature as 1.0). For other LLM evaluators, greedy search is used for reproducibility. We limit the computational budget to 71 evaluations on the validation dataset (including 21 evaluations during initiation in HPSS) for all prompt search methods (21 for Stepwise-Greedy as an exception, as their evaluation count is constant). Results of prompt search on Qwen2.5-14B-Instruct are averaged over 3 seeds and the standard deviation is provided. While for GPT-4o-mini, we report the results of one seed due to the API cost. More details including the cost of HPSS are in Appendix G.
4.2 Experimental Results
Human Alignment
The main results are presented in Table 2 and 3. Firstly, with the same generation times, HPSS substantially improves the performance of LLM evaluators compared to the starting point (i.e., the prompting strategy from MT-Bench), resulting in an average 29.4% relative improvement in Spearman correlation across various tasks and outperforming other Non-LLM and human-designed LLM evaluators by a large margin. Even with only 5% of the generation times, HPSS remains superior to G-Eval and CloserLook. Additional experiments in Appendix J also demonstrate that the Qwen2.5-14B-Instruct evaluator with HPSS substantially surpasses the human-designed Qwen2.5-72B-Instruct evaluator. Secondly, HPSS achieves a significant performance gain over other automatic prompt optimization methods, showing its better adaptability to NLG evaluation tasks. Finally, HPSS brings consistent performance improvements to different evaluation models, highlighting its cross-model generalization capability. We also validate that integrating HPSS with inference-time methods such as self-consistency Wang et al. (2023b) can further improve the performance of LLM evaluators in Appendix K.
Figure 3: Average performance of Qwen2.5-14B-Instruct evaluator under different validation dataset sizes on Summeval and Topical-Chat.
Ablation Study
In Table 2 and 3, we provide an ablation study on key components of HPSS for the Qwen-2.5-14B-Instruct evaluator, including the mutation type exploitation, the additional exploration term, and the entire heuristic-function-guided mutation mechanism. The results validate that all components contribute to the final performance of HPSS. Notably, removing the entire mechanism leads to the most significant performance degradation, highlighting its effectiveness. Moreover, the performance gains from the heuristic search mechanism vary across different datasets, probably due to different search difficulties.
The impact of validation dataset size and search budget.
Figure 3 shows the performance of HPSS under different validation dataset sizes. When validating with only 10% of the human expert annotations, the performance of HPSS remains marginally off and superior to human-designed LLM evaluators, demonstrating its effectiveness in low-resource scenarios. Figure 4 shows the performance of Qwen2.5-14B-Instruct evaluator under different search steps on Topical-Chat. While HPSS does not exhibit an advantage over baseline methods in the early stages of the search, it avoids getting trapped in local optima prematurely (which often appears in baseline methods) and ultimately converges into better solutions. We speculate that this can be attributed to two reasons: First, the mutation type exploitation expands the search scope of HPSS; Second, the advantage estimations for each value become more accurate as the search progresses, enabling an effective search even after good solutions have been obtained.
Figure 4: Performance of Qwen2.5-14B-Instruct evaluator under different search steps on Topical-Chat. The first 21 steps for HPSS are in Initialization, which also serves as the initial search history for OPRO, causing the two methods to produce identical results during these steps. Post-convergence results are omitted.
4.3 Analysis
The generalizability of prompting strategies across datasets.
Using Qwen2.5-14B-Instruct as the evaluation model, we directly apply the prompting strategies found by HPSS in Table 2 and 3 to other datasets. As shown in Table 4, these prompting strategies demonstrate strong generalizability across different datasets on the same evaluation aspects, as they still achieve significantly better evaluation performances than human-designed prompting strategies in most scenarios. We also observe that these prompting strategies show cross-aspect generalization of different datasets. However, the improvements in cross-aspect generalization are generally smaller than those in same-aspect generalization, compared to human-designed prompting strategies. The cross-dataset generalizability of prompting strategies may further reduce the overhead and improve the practicality of HPSS.
Aspects Source Target MT-Bench CloserLook HPSS
Dataset Dataset Prompt + ICL
Coherence Summeval HANNA 0.459 0.440 0.516
HANNA Summeval 0.449 0.478 0.505
Relevance Summeval HANNA 0.447 0.434 0.484
HANNA Summeval 0.451 0.480 0.556
Engagingness Topical-Chat HANNA 0.430 0.476 0.518
HANNA Topical-Chat 0.635 0.675 0.737
Naturalness Topical-Chat SFRES 0.311 0.396 0.411
SFRES Topical-Chat 0.497 0.580 0.566
Table 4: Performance of Qwen2.5-14B-Instruct evaluator when directly applying the prompting strategies found on the source datasets to the target datasets.
The advantages of different values.
We calculate the mean and standard deviations of the advantages for each factor at the end of HPSS across different datasets and evaluation aspects. Table 5 shows that the advantages exhibit considerable variance across tasks and evaluators, indicating the necessity of finding appropriate prompting strategies tailored to specific tasks and evaluators. However, there are some common characteristics: (1) A moderate scoring scale of 1-10 generally enhances evaluator performance, whereas a too coarse-grained scoring scale of 1-3 is less effective. (2) Under greedy decoding, directly generating scores and employing human-written evaluation criteria without AutoCoT, Metrics, and Reference generally improves evaluation performance. (3) Regarding the placement order of different components, positioning the task description at the beginning seems to yield better results, possibly because this arrangement is more logically coherent and helps evaluators concentrate more on sample information and evaluation criteria. These findings provide insights into future evaluation prompt design. We also provide a case study of HPSS in Appendix L.
Factors Values GPT-4o-mini Qwen2.5-14B
Scoring Scale 1-3 -3.545(1.828) -3.500(2.403)
1-5 -0.310(1.830) 0.393(1.792)
1-10 1.492(0.894) 1.967(1.652)
1-50 0.152(1.456) 0.462(1.259)
1-100 2.211(0.936) 0.678(1.994)
In-Context Example 0 -2.208(2.897) 0.487(3.363)
3 0.715(2.043) 0.860(3.193)
5 -0.049(2.826) 0.062(2.913)
10 1.542(3.012) -1.410(2.785)
Evaluation Criteria None -0.210(1.849) -0.457(1.936)
Human-Written 0.822(2.325) 0.283(2.158)
Self-Generated -0.612(1.885) 0.174(1.697)
Reference None 9.038(5.772) 5.513(4.732)
Self-Generated 3.968(6.404) -1.070(6.022)
Dialectic -13.00(11.62) -4.444(9.902)
Chain-of- Thought None 0.540(2.004) 0.318(1.475)
Prefix -0.560(2.033) -0.021(1.744)
Suffix 0.020(1.654) -0.297(1.277)
AutoCoT None 0.308(0.594) 0.346(1.989)
AutoCoT -0.308(0.594) -0.346(1.989)
Metrics None 2.191(2.575) 2.957(1.892)
Metrics -2.191(2.575) -2.957(1.892)
Order TD $\rightarrow$ ER $\rightarrow$ IC 1.128(1.617) 0.201(2.538)
TD $\rightarrow$ IC $\rightarrow$ ER 0.408(2.177) 1.240(2.376)
ER $\rightarrow$ TD $\rightarrow$ IC 0.711(1.970) -0.393(2.101)
ER $\rightarrow$ IC $\rightarrow$ TD 0.105(2.243) -0.506(3.965)
IC $\rightarrow$ TD $\rightarrow$ ER -0.449(2.254) 0.022(2.335)
IC $\rightarrow$ ER $\rightarrow$ TD -1.902(1.786) -0.565(2.836)
Table 5: Average advantages and standard deviations across different datasets and evaluation aspects for GPT-4o-mini and Qwen2.5-14B evaluators.
5 Conclusion
In this work, we integrate 8 key factors for the prompt of LLM evaluators and propose HPSS, a heuristic search method to automatically optimize the prompting strategy for adjusting these factors. HPSS leverages the expected performance advantages of each value as the heuristic function to guide the search process. Extensive experiments on four NLG evaluation tasks demonstrate the superiority of HPSS, yielding consistent performance gains over both human-designed evaluation prompts and automatic prompt optimization methods. Additionally, we analyze the characteristics of these factors and provide new insights into future prompt design for LLM evaluators.
6 Limitations
The limitations of our work are summarized as follows:
1) HPSS requires iterative search on a validation dataset with human annotations, introducing additional annotation and inference overheads. However, we believe that it is not a severe problem for the following three reasons: Firstly, the overall cost is affordable in the vast majority of scenarios. Under the default setting, where 50% of the dataset is selected as the validation dataset, the search cost for a specific evaluation aspect of one dataset is approximately \$8 for GPT-4o-mini, and approximately 40 minutes using 4 H100 GPUs for Qwen2.5-14B-Instruct (See Appendix G). Secondly, HPSS remains effective in low-resource scenarios. As shown in Figure 3, even when annotated data is scarce (reducing the validation dataset size to 10% of the entire dataset, $\sim$ 100 samples per dataset) , HPSS can still significantly outperform manually designed prompt templates. Finally, HPSS is efficient during inference on the test dataset, requiring only a single greedy decoding pass for each test sample while still achieving better performance than much larger manually designed LLM evaluators that require 20 generations and self-consistency decoding (See Table 2, Table 3, and Appendix J). Nonetheless, introducing well-designed early stopping strategies could potentially reduce the search overheads of HPSS, which is considered as important future work.
2) HPSS mainly focuses on the prompting strategy optimization of the input prompts for LLM evaluators. Despite the importance of input prompts, other modules such as decoding and interaction strategies may also benefit the performance of LLM evaluators. Regarding decoding strategies, our experiments demonstrate that incorporating self-consistency decoding can enhance the performance of HPSS (See Appendix K). Regarding multiple LLM evaluators’ interaction, in our preliminary experiments, we follow the settings of ChatEval Chan et al. (2024) to explore the impact of interaction strategies with multiple Qwen2.5-14B-Instruct evaluators, using the prompting strategies found by HPSS. We find that different interaction strategies have negligible impacts on the final evaluation performance, which is similar to the self-consistency method that directly averages the evaluation results of multiple individual LLM evaluators. Given that the search space increases significantly in multi-agent interaction scenarios and that it is difficult for LLM evaluators with different output formats to interact directly, we leave the optimization of interaction strategies among multiple LLM evaluators as important future work.
Appendix A List of Evaluation Prompt Templates
This section lists all prompt templates applied throughout this study, including the prompt templates utilized to generate the final rating (Table 9, 11, 13, 15) and the templates used to generate Reference, AutoCoT, and Metrics (Table 10, 12, 14, 16). For these prompt templates, we generally refer to the reference-free single answer pointwise grading prompt from MT-Bench Zheng et al. (2023). However, due to the contents of some components of this template being mixed, we make minor adjustments to the order of some sentences to ensure that the search for factor Order can be conducted. We use the prompt template of LLMBar Zeng et al. (2024) to generate Metrics.
Figure 5: Average dataset-level Spearman human correlation on Topical-Chat for GPT-4o-mini evaluator using different prompting strategies, which are modified based on the baseline strategy from MT-Bench for each factor.
Appendix B Preliminary Experiment on Factor Effects
To further explore the effect of each factor on the performance of LLM evaluators, we conducted a preliminary experiment with GPT-4o-mini on a commonly-used dialogue evaluation dataset Topical-Chat Gopalakrishnan et al. (2019). Using the prompting strategy from MT-Bench Zheng et al. (2023) with the default scoring scale 1-3 of Topical-Chat as the baseline, we adjust the selection value of each factor separately, employ greedy search decoding to generate rating scores, and report the Spearman correlations with human judgments. The results, as shown in Figure 5, indicate that these factors significantly influence the performance of LLM evaluators, yet some findings diverge from previous works. For instance, Pereira and Lotufo (2024) claims that Metrics can improve the performance of LLM evaluators. However, we observed a performance decrease on Topical-Chat. Chiang and Lee (2023) finds that CoT plays an important role in the evaluation prompt. However, when it comes to the GPT-4o-mini evaluator, adding CoT results in negligible differences. These results highlight the importance of optimizing the prompting strategy for adjusting these factors.
Appendix C Details of In-Context Example Selection
We perform stratified sampling based on human ratings within the validation dataset to obtain in-context examples, aiming to ensure an even distribution of examples across different human ratings. When evaluating the performance of some prompting strategies on the validation dataset, we remove the corresponding example in the in-context examples set if this example is to be evaluated, aiming to prevent data leakage.
Appendix D Details of Benchmarks
A brief introduction of the meta-evaluation benchmarks involved is as follows:
  • Summeval Fabbri et al. (2021) is a meta-evaluation benchmark for summarization. It contains human evaluation annotations for 16 summarization systems on 100 articles from the CNN / DailyMail corpus, resulting in a total of 1600 summary-level annotations. Each summary is evaluated on four aspects: Coherence, Consistency, Fluency, and Relevance. The authors recruit annotators on Amazon Mechanical Turk (AMT) to rate each summary on a scale from 1 to 5. Cross-validation by other annotators and experts is conducted to correct errors and enhance annotation quality.
  • Topical-Chat Gopalakrishnan et al. (2019) is a meta-evaluation benchmark for knowledge-grounded dialogue generation. It contains 360 samples, each including dialogue context, relevant knowledge, a response, and human ratings of the response across five aspects: Coherence, Engagingness, Groundedness, Naturalness, and Understandability, ranging from 1 to 3. The annotators recruited from AMT provide the ratings. Following HD-Eval Liu et al. (2024d), the first four aspects are used to measure the performance of HPSS.
  • SFHOT/ SFRES Wen et al. (2015) are meta-evaluation benchmarks for data-to-text generation. They contain 875 / 1181 samples respectively, which provide information about restaurants and hotels in San Francisco and aim to let the model generate corresponding utterances. The authors recruit annotators from AMT to rate the Informativeness and Naturalness of the generated utterances for each sample on a scale from 1 to 6.
  • HANNA Chhun et al. (2022) serves as a meta-evaluation benchmark for story generation. It contains 1,056 stories produced by 10 different automatic story generation systems. Each story is rated by 3 annotators recruited from Amazon Mechanical Turk on 6 aspects: Coherence, Relevance, Empathy, Surprise, Engagement, and Complexity. Ratings range from 1 to 5. The final score for each aspect is the average of the three annotators’ ratings.
  • MT-Bench Zheng et al. (2023) comprises 3.3k expert-level pairwise human evaluation of responses, generated by six LLMs on 80 carefully designed questions. These questions cover 7 categories: Writing, Roleplay, Reasoning Math, Coding, Extraction, STEM and Humanities. We select the first round of dialogues from this dataset and filter out the tied cases, leaving a final evaluation dataset of 1020 instances.
  • AUTO-J (Eval-P) Li et al. (2024) provides 1,392 pairwise comparison data, each of which contains a query, two LLM-generated responses, and a human-annotated preference label. This dataset involves 58 real-world scenarios and the responses are generated from 6 LLM families. We filter out the tied cases and leave a final evaluation dataset of 1019 instances.
  • LLMBar Zeng et al. (2024) is a meta-evaluation benchmark for instruction-following, which consists of two components: (1) The Natural set, which is gathered from existing human-preference datasets. (2) The Adversarial set, where the authors intentionally create misleading outputs that appear plausible but deviate from the instructions to challenge the evaluators. This dataset contains a total of 419 pairwise comparison instances.
Appendix E Metric Calculation
E.1 Pointwise Grading
Following previous work Liu et al. (2024d, c); Zhong et al. (2022), we adopt dataset-level (sample-level for Summeval as an exception) Spearman ($\rho$) correlation coefficient between human judgments and LLM evaluations to measure the performance of LLM evaluators. Given a dataset $\mathcal{D}$, evaluation aspect $a$ and evaluation metric $f(\cdot)$, we could calculate the human correlation of this evaluation metric at either dataset or sample level:
  • Dataset Level: For dataset-level human correlation, we evaluate the correlations on all samples in the dataset, as follows:
    $corr_{dataset}(\{s^{*}_{i,a}\}_{i=1}^{|D|},\{s_{i,a}\}_{i=1}^{|D|})=$ (7)
    $\rho([{s}^{*}_{i,a},...,{s}^{*}_{|D|,a}],[s_{i,a},...,s_{|D|,a}])$
    where $\{s^{*}_{i,a}\}_{i=1}^{|D|}$ and $\{s_{i,a}\}_{i=1}^{|D|}$ denote the evaluation results (for free-text evaluations, scores are extracted via rules as final evaluation results) for the aspect $a$ of dataset $\mathcal{D}$ from human annotations and evaluation metric $f(\cdot)$, respectively.
  • Sample Level: Assume that the dataset $\mathcal{D}$ consists of $J$ queries where each query has target responses from $M$ diverse systems (with a total of $|D|=M\times J$ samples), and for sample-level human correlation, we first compute correlations on multiple responses to an individual query (e.g., the summaries from 16 summarization systems on one article for Summeval), then average them across all queries:
    $corr_{sample}(\{s^{*}_{i,a}\}_{i=1}^{|D|},\{s_{i,a}\}_{i=1}^{|D|})=$ (8)
    $\frac{1}{J}\sum_{i=1}^{J}(\rho([{s}^{*}_{i1,a},...,{s}^{*}_{iM,a}],[s_{i1,a},...,s_{iM,a}]))$
    where $s^{*}_{ij,a}$ and $s_{ij,a}$ denote the evaluation results for the $j$-th response to $i$-th query for the aspect $a$ of dataset $\mathcal{D}$, from human annotations and evaluation metric $f(\cdot)$, respectively.
The evaluation metric $f(\cdot)$ is the LLM evaluator using a specific prompt template in our implementation, and the calculation for the Spearman correlation coefficient $\rho$ between two vectors of length $n$ is as follows:
$\rho=1-\frac{6\sum d_{i}^{2}}{n(n^{2}-1)}$ (9)
where $d_{i}$ represents the difference in the rank of the $i$-th element between two vectors, where the ranks are determined by sorting the elements within their respective vectors in ascending order.
E.2 Pairwise Comparison
Following previous work Zeng et al. (2024); Lambert et al. (2024), we directly adopt accuracy to measure the performance of LLM evaluators in pairwise comparison. Given a dataset $\mathcal{D}$, evaluation aspect $a$ and evaluation metric $f(\cdot)$, the accuracy is calculated as follows:
$acc(\{s^{*}_{i,a}\}_{i=1}^{|D|},\{s_{i,a}\}_{i=1}^{|D|})=\frac{1}{|D|}\mathbb{I}(s^{*}_{i,a}=s_{i,a})$ (10)
Appendix F Implementation Details of Baselines
As for APE, we use the LLM evaluator to resample new prompts. The queue size is set to 5. In each iteration, two new prompt candidates are resampled based on each prompt in the queue. As for OPRO, we use the LLM evaluator to generate new prompting strategies. We provide the LLM with the selection range for each factor, as well as the 20 previously top-performing strategies and their corresponding correlation metrics, which serve as the search history. The LLM is asked to generate a list containing the new selection strategy for each factor. The explored prompting strategies in Initiation will serve as the initial search history. As for Greedy, we perturb the current prompting strategy 5 times in each iteration by randomly replacing the value of one factor to generate new strategies. The strategy that performs best on the validation dataset is retained for the next iteration. Finally, as for Stepwise-Greedy, we follow the order shown in Table 1 to optimize each factor sequentially. In each step, we select the value for the current factor that performs best on the validation dataset while holding all other factors fixed. This selected choice is then established as the final optimization result for the current factor.
Appendix G Implementation Details of HPSS
We determine the hyperparameters for HPSS via grid search on the validation dataset of Topical-Chat using the Qwen2.5-14B-Instruct evaluator. Specifically, the population size $k$ is set to 5. The mutation time for each template $g=2$. The exploitation probability $\rho=0.2$. The temperature $\tau$ for the softmax function used to calculate the exploration probability of each template is set to 5, and the weight $\lambda$ for the additional exploration term is set to 4. We provide the performance of Qwen2.5-14B-Instruct after modifying each hyperparameter choice on the validation dataset of Topical-Chat in Figure 6.
Under the computational budget described in Section 4 (i.e., 71), the search cost of HPSS (i.e., the inference times of the LLM evaluator) for a specific evaluation aspect of one dataset is approximately 70 times the size of the validation dataset. When the size of the validation dataset is 50% of the entire dataset, for GPT-4o-mini, the cost of HPSS for a specific evaluation aspect of one dataset is approximately \$8. In total, the overall cost of HPSS across all the datasets and evaluation aspects is approximately \$140. For Qwen2.5-14B-Instruct, the runtime of HPSS for a specific evaluation aspect of one dataset is approximately 40 minutes using 4 H100 GPUs and the vllm Kwon et al. (2023) inference framework, while the overall runtime of HPSS across all the datasets and aspects is approximately 12 hours. The overall costs for other automatic prompt optimization methods for LLM evaluators are the same as HPSS, which is affordable in the vast majority of scenarios. We also validate that even with 1/5 of the above search cost (reducing the validation dataset size to 10% of the entire dataset), HPSS can still significantly outperform human-designed LLM evaluators in Figure 3.
Appendix H Details of Algorithm Implementation
We provide the detailed implementation of the two steps of HPSS in Algorithm 1 and 2 respectively.
Appendix I Experiments on Pairwise Comparison
I.1 Experimental Setup
Apart from pointwise grading tasks, we also validate our method on three pairwise comparison benchmarks, i.e., MT-Bench Zheng et al. (2023), AUTO-J Li et al. (2024), and LLMBar Zeng et al. (2024), which primarily focus on instruction-following tasks. Qwen2.5-14B-Instruct is employed as the evaluation model, with all hyperparameters remaining the same as pointwise grading experiments. Regarding the search space, we remove the factor Scoring Scale, and the value Self-Generated Criteria of the factor Evaluation Criteria, which only exist in the pointwise grading setting. We compare our method with the prompting strategies from MT-Bench, the best human-designed prompting strategies Metrics+Reference found by LLMBar, and all automatic prompt optimization methods examined in pointwise grading experiments. Results of prompt search are averaged over 3 random seeds and the standard deviation is provided.
I.2 Main Results
The main results are provided in Table 6. HPSS substantially improves the performance of LLM evaluators compared to human-designed prompting strategies and achieves the best average performance across all automatic prompt optimization methods, which validates its effectiveness in pairwise comparison prompt optimization. On the AUTO-J (Eval-P), multiple prompt optimization methods simultaneously achieve the best results. Upon data examination, we find that the best prompting strategy is close to the starting point, with only one factor having a different value, resulting in lower search difficulty. Overall, we observe that human-designed prompting strategies for pairwise comparison tasks have already been well-optimized, and the improvements brought by automatic prompt optimization are relatively modest compared to those in pointwise grading tasks.
Method AUTO-J LLMBar MT-Bench Avg.
(Eval-P)
MT-Bench 0.792 0.619 0.765 0.725
Metrics+Reference* 0.800 0.724 0.778 0.767
APE 0.799(0.005) 0.670(0.032) 0.774(0.008) 0.748
OPRO 0.847(0.000) 0.695(0.000) 0.791(0.005) 0.778
Greedy 0.820(0.019) 0.774(0.005) 0.775(0.013) 0.790
Stepwise-Greedy 0.847(0.000) 0.743(0.000) 0.784(0.000) 0.791
HPSS (Ours) 0.847(0.000) 0.778(0.005) 0.789(0.015) 0.805
Table 6: Accuracy of different prompting methods based on Qwen-2.5-14B-Instruct on pairwise comparison datasets including AUTO-J (Eval-P), LLMBar, and MT-Bench.
Model Method Summeval Topical-Chat SFHOT HANNA
Qwen-2.5-72B-Instruct MT-Bench 0.502 0.634 0.362 0.460
CloseLook + ICL 0.481 0.592 0.316 0.470
CloseLook + ICL 0.525 0.623 0.300 0.517
Qwen-2.5-14B-Instruct HPSS 0.560 0.697 0.431 0.519
Table 7: Average performance comparison of Qwen2.5-72B-Instruct evaluator and Qwen2.5-14B-Instruct evaluator on Summeval, Topical-Chat, SFHOT, and HANNA. †  indicates that the corresponding method employs 20 generations with self-consistency.
Method Summeval Topical-Chat SFHOT HANNA
CloseLook + ICL 0.532 0.660 0.402 0.555
HPSS 0.564 0.740 0.431 0.545
HPSS 0.569 0.763 0.453 0.578
Table 8: Average performance of HPSS under different generation times on Summeval, Topical-Chat, SFHOT, and HANNA. †  indicates that the corresponding method employs 20 generations with self-consistency. GPT-4o-mini is employed as the evaluation model.
Figure 6: Performance of Qwen2.5-14B-Instruct evaluator under different hyperparameters settings. We provide the average results over 3 seeds on the validation dataset of Topical-Chat.
Appendix J Comparison with Stronger LLM Evaluator
As shown in Table 7, we compare the performance of the Qwen2.5-14B-Instruct evaluator using the prompting strategies obtained by HPSS with the Qwen2.5-72B-Instruct evaluator using human-designed prompting strategies. Notably, Qwen2.5-14B-Instruct with HPSS achieve significantly better evaluation performance than the human-designed Qwen2.5-72B-Instruct evaluator, even with only 5% of the generation times. These results demonstrate the efficiency of HPSS.
Appendix K Incorporating HPSS with Inference-Time Methods
As illustrated in Table 3, for GPT-4o-mini evaluator, the performance of CloserLook + ICL surpasses HPSS on HANNA. Considering the different inference overheads of these two methods, we attempt to incorporate self-consistency Wang et al. (2023b) decoding strategy into HPSS to ensure a fair comparison. Specifically, we conduct 20 generations and compute the average evaluation score, which is consistent with the setting of CloserLook + ICL. As shown in Table 8, self-consistency further enhances the performance of HPSS, consistently outperforming CloserLook + ICL with the same generation times. This result indicates that integrating HPSS with inference-time methods can further improve the performance of LLM evaluators.
Algorithm 1 Initialization
1:$n$ factors $F_{1},F_{2},\ldots,F_{n}$, baseline prompting strategy $\mathcal{T}^{base}=\mathcal{T}_{f_{1b_{1}},f_{2b_{2}},\ldots,{f_{nb_{n}}}}$, performance metrics $r$
2:$T\leftarrow\{\mathcal{T}^{base}\}$
3:$s_{base}\leftarrow c(\mathcal{T}^{base})$
4:for $i=1,\cdots,n$ do
5:     for $j=1,\cdots,m_{i}\ \text{and}\ j\neq b_{i}$ do
6:         $T\leftarrow T\cup\{\mathcal{T}_{f_{1b_{1}},f_{2b_{2}},\ldots,f_{ij},\ldots,{f_{nb_{n}}}}\}$
7:         $s_{ij}\leftarrow r(\mathcal{T}_{f_{1b_{1}},f_{2b_{2}},\ldots,f_{ij},\ldots,{f_{nb_{n}}}})$
8:     end for
9:     for $j=1,\cdots,m_{i}$ do
10:         $A_{ij}\leftarrow s_{ij}-avg(s_{i1},s_{i2},\cdots,s_{i{m_{i}}})$
11:     end for
12:end for
13:return the best $k$ prompt templates based on $s$, as $best$
Algorithm 2 Iterative Search
1:Explored prompting strategies set $T$, initial prompting strategies population $best$, performance metrics $r$, beam size $k$, budget $m$, hyper-parameters $\lambda,g,\tau,\rho$
2:$cost\leftarrow 0$
3:while $cost<m$ do
4:     $new\_best\leftarrow best$
5:     for each $\mathcal{T}^{c}$ in $best$ do
6:         for $l=1,\cdots,g$ do
7:              Sample new candidate prompting strategy $\mathcal{T}^{new}$ in $T_{adj}$ based on Equation 3
8:              if $\mathcal{T}^{new}\in T$ then
9:                  continue $\triangleright$ Skip explored strategies
10:              end if
11:              Sample $\alpha\sim\text{Bernoulli}(\rho)$
12:              if $\alpha$ then
13:                  Select $\mathcal{T}^{max}$ with the highest overall advantage and $\mathcal{T}^{max}\notin T$ based on Equation 6
14:                  $\mathcal{T}^{new}\leftarrow\mathcal{T}^{max}$
15:              end if
16:              $T\leftarrow T\cup\{\mathcal{T}^{new}\}$
17:              $s^{new}\leftarrow r(\mathcal{T}^{new})$
18:              $new\_best\leftarrow new\_best\cup\{\mathcal{T}_{new}\}$
19:              $cost\leftarrow cost+1$
20:              if not $\alpha$ then
21:                  Update advantage $A$ based on Equation 4 and 5
22:              end if
23:         end for
24:     end for
25:     Select the best $k$ prompting strategies in $new\_best$, as $best$
26:end while
27:return $best[0]$
Template Value Prompt
Backbone -
## Instruction
Please act as an impartial judge and evaluate the quality of the summary of the news article displayed below on its [Aspect]. {reference_1_template} {reference_dialectic_template} {chain_of_thought_template}
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the summary. [Criteria]
2. Be as objective as possible.
{autocot_template}
{in_context_example_template}
## Article
[Article]
{metrics_template}
{reference_2_template}
## The Start of the Summary
[Summary]
## The End of the Summary
Reference 1 - You will be given the news article, the summary, and a high-quality reference summary.
Reference 2 - ## The Start of Reference Summary
[Reference]
## The End of Reference Summary
Reference Dialectic - Please generate your own summary for the news article first and take into account your own summary to evaluate the quality of the given summary.
Chain-of-Thought No CoT You must directly output your rating of the summary on a scale of 1 to {max} without any explanation by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Prefix CoT Begin your evaluation by providing a short explanation. After providing your explanation, you must rate the summary on a scale of 1 to {max} by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Suffix CoT You must rate the summary on a scale of 1 to {max} first by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]". And then provide your explanation.
Scoring Scale 3 {max} $=3$
5 {max} $=5$
10 {max} $=10$
50 {max} $=50$
100 {max} $=100$
AutoCoT - Evaluation Steps:
[Autocot]
In-Context Example - Here are some examples and their corresponding ratings:
{example_template_1}
{example_template_2}
{example_template_n}
Following these examples, evaluate the quality of the summary of the news article displayed below on its [Aspect]:
Example - ## Example [Number]:
## Article
[Article]
## The Start of the Summary
[Summary]
## The End of the Summary
## Rating
[Human Rating]
Metrics - ## Questions about Summary
Here are some questions about the summary. You can do the evaluation based on thinking about all the questions.
[Metrics]
Table 9: Detailed evaluation prompt templates for Summeval. The backbone serves as the final input prompt template for LLM evaluators. The three components marked in different colors represent Task Description (TD), Evaluation Rule (ER), and Input Content (IC), respectively. The content within {} represents the prompt template for each factor, corresponding to the following rows in this table. Different content may be chosen for each template when corresponding factor values vary. The content within [] is sample-specific input information. - in Value means that when the factor is chosen as "None", this template will be replaced with an empty string (""). Otherwise, the content of this template will be added to the backbone. Specifically, the templates Reference 1 and Reference 2 will be replaced with an empty string ("") unless the factor Reference is chosen as Self-Generated Reference. The template Reference Dialectic will be replaced with an empty string ("") unless the factor Reference is chosen as Dialectic.
Template Prompt
Reference Generation Please summarize the following text: [Article]
Summary:
AutoCoT Generation ## Instruction
Please act as an impartial judge and evaluate the quality of the summary of the news article on its [Aspect] and rate the summary on a scale of 1 to {max}.
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the summary. [Criteria]
2. Be as objective as possible.
Please generate the evaluation steps for this task without other explanation.
Evaluation Steps:
Metrics Generation ## Instruction
Please act as an impartial judge and evaluate the quality of the summary of the news article displayed below on its [Aspect]. Please propose at most three concise questions about whether a potential summary is a good summary for a given news article on its [Aspect]. Another assistant will evaluate the aspect of the summary by answering all the questions.
Here are some rules of the evaluation:
(1) Your evaluation should consider the [Aspect] of the summary. [Criteria]
(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
## Article:
[Article]
## Requirements for Your Output:
(1) The questions should **specifically** target the given news article instead of some general standards so that the questions may revolve around key points of the news article.
(2) You should directly give the questions without any other words.
(3) Questions are presented from most important to least important.
Table 10: Detailed prompt templates for Reference, AutoCoT, and Metrics generation for Summeval.
Template Value Prompt
Backbone -
## Instruction
Please act as an impartial judge and evaluate the quality of the response for the next turn in the conversation displayed below on its [Aspect]. The response concerns an interesting fact, which will be provided as well. {reference_1_template} {reference_dialectic_template} {chain_of_thought_template}
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the response. [Criteria]
2. Be as objective as possible.
{autocot_template}
{in_context_example_template}
## Conversation History
[Conversation History]
{metrics_template}
## Corresponding Fact
[Corresponding Fact]
{reference_2_template}
## The Start of Response
[Response]
## The End of the Response
Reference 1 - You will also be given a high-quality reference response with the conversation.
Reference 2 - ## The Start of Reference Response
[Reference]
## The End of Reference Response
Reference Dialectic - Please generate your own response for the next turn in the conversation first and take into account your own response to evaluate the quality of the given response.
Chain-of-Thought No CoT You must directly output your rating of the response on a scale of 1 to {max} without any explanation by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Prefix CoT Begin your evaluation by providing a short explanation. After providing your explanation, you must rate the response on a scale of 1 to {max} by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Suffix CoT You must rate the response on a scale of 1 to {max} first by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]". And then provide your explanation.
Scoring Scale 3 {max} $=3$
5 {max} $=5$
10 {max} $=10$
50 {max} $=50$
100 {max} $=100$
AutoCoT - Evaluation Steps:
[Autocot]
In-Context Example - Here are some examples and their corresponding ratings:
{example_template_1}
{example_template_2}
{example_template_n}
Following these examples, evaluate the quality of the response for the next turn in the conversation displayed below on its [Aspect]:
Example - ## Example [Number]:
## Conversation History
[Conversation History]
## Corresponding Fact
[Corresponding Fact]
## The Start of the Response
[Response]
## The End of the Response
## Rating
[Human Rating]
Metrics - ## Questions about Response
Here are some questions about the response. You can do the evaluation based on thinking about all the questions.
[Metrics]
Table 11: Detailed evaluation prompt templates for Topical-Chat.
Template Prompt
Reference Generation Please output the response for the next turn in the conversation. Conversation History: [Conversation History]
Response:
AutoCoT Generation ## Instruction
Please act as an impartial judge and evaluate the quality of the response for the next turn in the conversation on its [Aspect] and rate the response on a scale of 1 to {max}.
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the response. [Criteria]
2. Be as objective as possible.
Please generate the evaluation steps for this task without other explanation.
Evaluation Steps:
Metrics Generation ## Instruction
Please act as an impartial judge and evaluate the quality of the response for the next turn in the conversation displayed below on its [Aspect]. Please propose at most three concise questions about whether a potential response is a good response for the next turn in the given conversation on its [Aspect]. Another assistant will evaluate the aspect of the output by answering all the questions.
Here are some rules of the evaluation:
(1) Your evaluation should consider the [Aspect] of the response. [Criteria]
(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
## Conversation History:
[Conversation History]
## Requirements for Your Output:
(1) The questions should **specifically** target the given conversation instead of some general standards, so the questions may revolve around key points of the conversation.
(2) You should directly give the questions without any other words.
(3) Questions are presented from most important to least important.
Table 12: Detailed prompt templates for Reference, AutoCoT, and Metrics generation for Topical-Chat.
Template Value Prompt
Backbone -
## Instruction
Please act as an impartial judge and evaluate the quality of a natural language sentence generated according to a structured data expression displayed below on its [Aspect]. {reference_1_template} {reference_dialectic_template} {chain_of_thought_template}
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the sentence. [Criteria]
2. Be as objective as possible.
{autocot_template}
{in_context_example_template}
## Structured Data Expression
[Structured Data Expression]
{metrics_template}
{reference_2_template}
## The Start of the Natural Language Sentence
[Natural Language Sentence]
## The End of the Natural Language Sentence
Reference 1 - You will be given the structured data expression, the sentence and a high-quality reference sentence.
Reference 2 - ## The Start of Reference Sentence
[Reference]
## The End of Reference Sentence
Reference Dialectic - Please generate your own sentence according to the given structured data expression first and take into account your own sentence to evaluate the quality of the given sentence.
Chain-of-Thought No CoT You must directly output your rating of the sentence on a scale of 1 to {max} without any explanation by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Prefix CoT Begin your evaluation by providing a short explanation. After providing your explanation, you must rate the sentence on a scale of 1 to {max} by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Suffix CoT You must rate the sentence on a scale of 1 to {max} first by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]". And then provide your explanation.
Scoring Scale 3 {max} $=3$
5 {max} $=5$
10 {max} $=10$
50 {max} $=50$
100 {max} $=100$
AutoCoT - Evaluation Steps:
[Autocot]
In-Context Example - Here are some examples and their corresponding ratings:
{example_template_1}
{example_template_2}
{example_template_n}
Following these examples, evaluate the quality of a natural language sentence generated according to a structured data expression displayed below on its [Aspect]:
Example - ## Example [Number]:
## Structured Data Experssion
[Structured Data Experssion]
## The Start of the Natural Language Sentence
[Natural Language Sentence]
## The End of the Natural Language Sentence
## Rating
[Human Rating]
Metrics - ## Questions about Sentence
Here are some questions about the sentence. You can do the evaluation based on thinking about all the questions.
[Metrics]
Table 13: Detailed evaluation prompt templates for SFHOT / SFRES.
Template Prompt
Reference Generation Please generate a natural language sentence generated according to a structured data expression. Expression: [Expression]
Sentence:
AutoCoT Generation ## Instruction
Please act as an impartial judge and evaluate the quality of a natural language sentence generated according to a structured data expression displayed below on its [Aspect] and rate the sentence on a scale of 1 to {max}.
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the sentence. [Criteria]
2. Be as objective as possible.
Please generate the evaluation steps for this task without other explanation.
Evaluation Steps:
Metrics Generation ## Instruction
Please act as an impartial judge and evaluate the quality of a natural language sentence generated according to a structured data expression displayed below on its [Aspect]. Please propose at most three concise questions about whether a potential sentence is a good sentence generated according to a given structured data expression on its [Aspect]. Another assistant will evaluate the aspects of the sentence by answering all the questions.
Here are some rules of the evaluation:
(1) Your evaluation should consider the [Aspect] of the sentence. [Criteria]
(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
## Structured Data Expression:
[Structured Data Expression]
## Requirements for Your Output:
(1) The questions should **specifically** target the given structured data expression instead of some general standards, so the questions may revolve around key points of the structured data expressions.
(2) You should directly give the questions without any other words.
(3) Questions are presented from most important to least important.
Table 14: Detailed prompt templates for Reference, AutoCoT, and Metrics generation for SFHOT / SFRES.
Template Value Prompt
Backbone -
## Instruction
Please act as an impartial judge and evaluate the quality of the story generated according to a prompt displayed below on its [Aspect]. {reference_1_template} {reference_dialectic_template} {chain_of_thought_template}
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the story. [Criteria]
2. Be as objective as possible.
{autocot_template}
{in_context_example_template}
## Prompt
[Prompt]
{metrics_template}
{reference_2_template}
## The Start of the Story
[Story]
## The End of the Story
Reference 1 - You will be given the prompt, the generated story and a high-quality reference story.
Reference 2 - ## The Start of Reference Story
[Reference]
## The End of Reference Story
Reference Dialectic - Please generate your own story for the given prompt first and take into account your own story to evaluate the quality of the given story.
Chain-of-Thought No CoT You must directly output your rating of the story on a scale of 1 to {max} without any explanation by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Prefix CoT Begin your evaluation by providing a short explanation. After providing your explanation, you must rate the story on a scale of 1 to {max} by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]".
Suffix CoT You must rate the story on a scale of 1 to {max} first by strictly following this format: "[[rating]]", for example: "Rating: [[{max}]]". And then provide your explanation.
Scoring Scale 3 {max} $=3$
5 {max} $=5$
10 {max} $=10$
50 {max} $=50$
100 {max} $=100$
AutoCoT - Evaluation Steps:
[Autocot]
In-Context Example - Here are some examples and their corresponding ratings:
{example_template_1}
{example_template_2}
{example_template_n}
Following these examples, evaluate the story generated according to a prompt displayed below on its [Aspect]:
Example - ## Example [Number]:
## Prompt
[Prompt]
## The Start of the Story
[Story]
## The End of the Story
## Rating
[Human Rating]
Metrics - ## Questions about Story
Here are some questions about the story. You can do the evaluation based on thinking about all the questions.
[Metrics]
Table 15: Detailed evaluation prompt templates for HANNA.
Template Prompt
Reference Generation Please generate a story according to the given prompt: [Prompt]
Story:
AutoCoT Generation ## Instruction
Please act as an impartial judge and evaluate the quality of the story generated according to a prompt displayed below on its [Aspect] and rate the story on a scale of 1 to {max}.
Here are some rules of the evaluation:
1. Your evaluation should consider the [Aspect] of the story. [Criteria]
2. Be as objective as possible.
Please generate the evaluation steps for this task without other explanation.
Evaluation Steps:
Metrics Generation ## Instruction
Please act as an impartial judge and evaluate the quality of the story generated according to a prompt displayed below on its [Aspect]. Please propose at most three concise questions about whether a potential story is a good story according to a given prompt on its [Aspect]. Another assistant will evaluate the aspect of the story by answering all the questions.
Here are some rules of the evaluation:
(1) Your evaluation should consider the [Aspect] of the story. [Criteria]
(2) Outputs should NOT contain more/less than what the instruction asks for, as such outputs do NOT precisely execute the instruction.
## Prompt:
[Prompt]
## Requirements for Your Output:
(1) The questions should **specifically** target the given prompt instead of some general standards, so the questions may revolve around key points of the prompt.
(2) You should directly give the questions without any other words.
(3) Questions are presented from most important to least important.
Table 16: Detailed prompt templates for Reference, AutoCoT, and Metrics generation for HANNA.
Model Dataset Aspect During Prompting Strategy Spearman
HPSS Scale Example Criteria Reference CoT AutoCoT Metrics Order Correlation
GPT-4o-mini SFHOT Naturalness Origin 5 0 Human Prefix TD $\rightarrow$ ER $\rightarrow$ IC 0.366
Best 100 0 Human TD $\rightarrow$ ER $\rightarrow$ IC 0.437 (+19.4%)
Worst 5 0 Human Dialectic Prefix TD $\rightarrow$ ER $\rightarrow$ IC 0.104 (-71.6%)
HANNA Complexity Origin 5 0 Human Prefix TD $\rightarrow$ ER $\rightarrow$ IC 0.458
Best 100 10 Self-Generated IC $\rightarrow$ ER $\rightarrow$ TD 0.617 (+34.7%)
Worst 5 0 Human Dialectic Prefix TD $\rightarrow$ ER $\rightarrow$ IC -0.046 (-110.0%)
Qwen2.5-14B SFRES Informativeness Origin 5 0 Human Prefix TD $\rightarrow$ ER $\rightarrow$ IC 0.169
Best 50 3 Dialectic Prefix TD $\rightarrow$ IC $\rightarrow$ ER 0.381 (+125.4%)
Worst 5 5 Prefix ER $\rightarrow$ IC $\rightarrow$ TD 0.088 (-47.9%)
Topical-Chat Coherence Origin 3 0 Human Prefix TD $\rightarrow$ ER $\rightarrow$ IC 0.520
Best 10 0 Self-Generated Suffix ER $\rightarrow$ IC $\rightarrow$ TD 0.651 (+25.2%)
Worst 10 5 Dialectic ER $\rightarrow$ TD $\rightarrow$ IC 0.456 (-12.3%)
Table 17: Illustration of some prompting strategies explored by HPSS and their performance on the validation dataset. We present both the best-performing and worst-performing prompting strategies for specific tasks during HPSS.
Appendix L Case Study
In Table 17, we present some cases of prompting strategies explored by HPSS for specific tasks. We find that there is a significant performance gap between the best-performing and worst-performing prompting strategies, emphasizing the sensitivity of LLM evaluators to prompting strategy. Figure 7 and 8 illustrate the original evaluation prompts from MT-Bench alongside the optimized one found by HPSS, specifically focusing on the aspect Complexity in HANNA and the aspect Coherence in Topical-Chat. The prompting strategies of these prompts are shown in the "origin" and "best" rows in Table 17 for their respective datasets and aspects. We observe that some of the prompting strategies found by HPSS include values that are rarely considered in human-designed evaluation prompts. For instance, for the aspect Complexity in HANNA, the prompting strategy found by HPSS places the input content (IC) first and the task description (TD) last. For the aspect Coherence in Topical-Chat, evaluation criteria are not used in HPSS. These results demonstrate the limitations of manual prompt design and underscores the importance of automatic prompting strategy optimization.
Furthermore, to intuitively show the reason why HPSS achieves better evaluation performance than human-designed LLM evaluators, we provide two judgment generation cases for different LLM evaluators in Table 18 and 19. The corresponding evaluation prompts are shown in Figure 9 and 10. In the first case from the text summarization task presented in Table 18, HPSS provides a balanced assessment by analyzing both strengths and weaknesses of the summary, with emphasis on overall fluency. In contrast, MT-Bench and CloserLook + ICL overemphasize minor issues like typos while overlooking its advantages in overall fluency, and exhibit some hallucinations in judgment. In the second case from the story generation task presented in Table 19, HPSS conducts a systematic analysis of the story and correctly identifies both key plots that effectively convey emotions and overly idealized plots that weaken emotional delivery. In contrast, MT-Bench and CloserLook + ICL each overlook one of these two important points, resulting in less accurate evaluations. These observations demonstrate that HPSS improves the ability of LLM evaluators to conduct comprehensive evaluations and achieve a balanced assessment of strengths and weaknesses within the input sample.
Figure 7: The original evaluation prompt for the aspect Complexity in HANNA from MT-Bench and the corresponding evaluation prompt found by HPSS for GPT-4o-mini evaluator. Factors with values other than "None" in the evaluation prompts are highlighted. The three main components of the evaluation prompt are annotated on the right side, e.g., Task Description (TD), Evaluation Rule (ER), and Input Content (IC). The placement order of these three components is also considered a factor in our optimization.
Figure 8: The original evaluation prompt for the aspect Coherence in Topical-Chat from MT-Bench and the corresponding evaluation prompt found by HPSS for Qwen2.5-14B-Instruct evaluator.
Figure 9: The original evaluation prompt for the aspect Fluency in Summeval from MT-Bench and the corresponding evaluation prompt found by HPSS for Qwen2.5-14B-Instruct evaluator.
Figure 10: The original evaluation prompt for the aspect Empathy in HANNA from MT-Bench and the corresponding evaluation prompt found by HPSS for Qwen2.5-14B-Instruct evaluator.
Article Chelsea have made an offer for FC Tokyo’s 22-year-old forward Yoshinori Muto, according to club president Naoki Ogane.
The Japan international, who has played for the J-League side since 2013, will join Chelsea’s Dutch partner club Vitesse Arnhem on loan next season if he completes a move to Stamford Bridge this summer.
Ogane claims that Chelsea’s interest in Muto is not connected to the £200million sponsorship deal they signed with Japanese company Yokohama Rubber in February.
FC Tokyo forward Yoshinori Muto (centre) brings the ball forward against Albirex Niigata in March.
FC Tokyo president Naoki Ogane claims that Chelsea have made a bid for Japan international Muto.
Muto tussles with Yuji Nakazawa of Yokohama F.Marinos during a J-League clash last month.
YOSHINORI MUTO FACTFILE
Age: 22
Club: FC Tokyo
Appearances: 37
Goals: 16
International caps (Japan): 11
International goals: 1
Did you know? Muto graduated from Keio University in Tokyo with an economics degree two weeks ago.
Speaking to Sports Nippon, Ogane said: ’It is true that Chelsea sent us an offer for Muto. ’It is a formal offer with conditions. They want to acquire him in the summer.’
Muto, who only graduated from Keio University a fortnight ago after completing an economics degree, would be the first Japanese player to represent Chelsea if he moves to west London. He has earned 11 caps for his country after signing his first professional contract in 2014, scoring once for the Samurai Blue.
A £4million deal for the youngster has been mooted, but Muto admits that he isn’t sure if he will join the Premier League title chasers despite being pleased with their bid.
He said: ’I have not decided yet at all. It is an honour for me to receive this offer from a great club.’
Muto scored 13 times in his debut season with FC Tokyo and was named in the J-League’s best XI.
Muto admits it is an ’honour’ to receive an offer from Chelsea although he has not yet decided to join the club.
Muto, pictured in action against Ventforet Kofu has scored three goals in four games so far this season.
The 22-year-old has a shot at goal during Japan’s Asian Cup match against Palestine in January this year.
He has continued his fine form during the current campaign, helping his club to third place in the division with three goals in four games.
Yokohama Rubber - one of the world’s largest tyre manufacturers - will become Chelsea’s official shirt sponsors from the start of the 2015-16 season.
The initial five-year deal is the biggest in the club’s history, with the Blues now considering a two-week pre-season tour of Japan this summer.
Summary Chelsea have made an offer for fc tokyo’s 22-year-old forward yoshinori muto. The japan international will join chelsea’s dutch partner club vitesse arnhem on loan next season if he completes a move to stamford bridge this summer. Muto claims chelsea’s interest in muto is not connected to the # 200million sponsorship deal they signed with japanese company yokohama rubber in february.
Evaluation Aspect Fluency
Human Score (1-5) 5
MT-Bench (1-5) The summary contains several issues related to fluency, primarily due to grammatical errors, inconsistencies in capitalization, and the inclusion of a symbol (#) that does not belong. The sentences are generally clear but lack the smoothness and refinement that one would expect in a quality summary. The use of lower-case letters for "fc tokyo" and "stamford bridge" is incorrect, and the number symbol (#) before "200million" is out of place. There are also minor issues with sentence structure and flow, such as the abrupt transition from the offer to the loan arrangement without proper context.
Rating: [[2]]
CloserLook + ICL (1-5) Analysis: The summary contains several grammatical errors and lacks coherence. Sentence fragments and improper punctuation are present, making the summary difficult to read. For example, "The japan international will join chelsea’s vitesse arnhem on loan next season if he completes a move to stamford bridge this summer" is not a complete sentence and lacks clarity. Additionally, the summary includes a hashtag (# 200million) instead of writing out the number, which is inconsistent with standard writing practices. The summary also omits some important details from the original article, such as Muto’s achievements and his indecision about the move.
Rating: 1
HPSS (1-10) Rating: [[8]]
Explanation:
The summary is generally well-written and grammatically correct. The sentences flow well and are coherent, providing a clear and concise overview of the key points from the article. However, there are a few areas where the fluency could be improved:
1. The summary starts with a run-on sentence that combines two separate ideas from the article: "Chelsea have made an offer for fc tokyo’s 22-year-old forward yoshinori muto. The japan international will join chelsea’s vitesse arnhem on loan next season if he completes a move to stamford bridge this summer." This could be split into two sentences for better readability: "Chelsea have made an offer for FC Tokyo’s 22-year-old forward Yoshinori Muto. If he completes a move to Stamford Bridge this summer, the Japan international will join Chelsea’s Vitesse Arnhem on loan next season."
2. The use of "# 200million" instead of "200 million" is inconsistent with the currency used in the original article and can be confusing for readers.
3. While the summary is concise, it could include more details to better reflect the depth of the original article. For instance, mentioning Muto’s form this season or his academic background could enhance the summary without significantly increasing its length.
Overall, the summary is clear and accurately reflects the main points of the article, but a few minor adjustments could improve its fluency and completeness.
Table 18: A judgment generation case of Qwen2.5-14B-Instruct evaluator using the prompting strategy from MT-Bench, CloserLook + ICL, and HPSS for the aspect Fluency in Summeval. Factual errors in the judgments are highlighted in red.
Prompt One normal day you see a mouse pointer move across your field of vision, right click, and delete an object.
Story It was a long tiring day as usual. I leaned on my sofa throwing my bag on to the near by table. Slowly dozed off to sleep. My imaginations rushed in, I could literally feel my dream and it was of none other than an episode which caught my attention today. Me and one of my colleague were having lunch and a small boy standing outside the canteen was cleaning the window panes, he was hardly 10. Full greased head to foot, filthy clothes, torn shoes but with a cherry red lips and a droopy eyes filled with all innocence he could gather within it. It was raining and he was shivering after a while. He sat loopy near the window base facing me. I was having coffee then. As we started staring at each other I felt difficulty in drinking my coffee. It was nothing but me sitting comfortably inside the cafe warm and the boy now half drenched killed me within. I could n’t read any thing from his face that he is not liking what he is going through. I called him and offered him a tea. Gladly he accepted it! I asked dont you feel cold, he told no I feel close to nature. I was so taken by the love and spirit he had within and his love for nature, the smallest things which I could n’t realize which he made me feel was aw inspiring. I wish I had a mouse with me to take of the poverty in his life with a click, and fill his childhood with all the beautiful things he deserve. And soon i realized this was a beautiful dream, a dream that i would love it to come true. Daffodils
Evaluation Aspect Empathy
Human Score (1-5) 3.67
MT-Bench (1-5) The story does a remarkable job of evoking empathy by deeply immersing the reader in the narrator’s emotional journey. The narrator’s guilt and compassion are palpable as they reflect on the stark contrast between their comfortable existence and the boy’s challenging circumstances. The emotional resonance is further heightened by the narrator’s wish to alleviate the boy’s poverty with a simple click, symbolizing a desire to change the world through a small, symbolic action. The dream sequence provides a poignant backdrop that amplifies the narrator’s sentiments.
Rating: [[5]]
CloserLook + ICL (1-5) Analysis: The story does not directly address the prompt about a mouse pointer deleting an object. Instead, it focuses on a dream about a young boy cleaning window panes in the rain, highlighting the character’s emotional response to the boy’s innocence and resilience. The narrator feels a deep sense of empathy and guilt, wishing they could alleviate the boy’s poverty. While the story does evoke strong emotions, it does not connect these emotions to the prompt’s scenario involving a mouse pointer.
Rating: 3
HPSS (1-100) The story delves into the narrator’s emotional response to a young boy cleaning window panes outside a canteen during a rainy day. The narrator’s empathy is palpable as they describe the stark contrast between their comfortable situation inside the cafe and the boy’s difficult circumstances. The emotions conveyed include compassion, guilt, and a strong desire to alleviate the boy’s poverty. The narrator’s emotional journey from discomfort to connection is well-articulated and relatable. However, the narrative occasionally veers into a dream sequence that includes the wish for a magical solution to poverty, which slightly detracts from the realism and depth of the emotional engagement. Despite this, the core emotional experience is effectively conveyed and deeply felt.
Rating: [[75]]
Table 19: A judgment generation case of Qwen2.5-14B-Instruct evaluator using the prompting strategy from MT-Bench, CloserLook + ICL, and HPSS for the aspect Empathy in HANNA.